A study of Hub Furniture regarding the payment of invoices reveals the time from billing until payment is received follows the normal distribution. The mean time until payment is received is 20 days and the standard deviation is 5 days.
a. What percent of the invoices are paid within 15 days of receipt?
b. What percent of the invoices are paid in more than 28 days?
c. What percent of the invoices are paid in more than 15 days but less than 28 days?
d. The management of Hub Furniture wants to encourage their customers to pay their monthly invoices as soon as possible. Therefore, it announced that a 2 percent reduction in price would be in effect for customers who pay within 7 days of the receipt of the invoice. What percent of customers will earn this discount?

Answer by stanbon(48568)

A study of Hub Furniture regarding the payment of invoices reveals the time from billing until payment is received follows the normal distribution. The mean time until payment is received is 20 days and the standard deviation is 5 days.
a. What percent of the invoices are paid within 15 days of receipt?
z(15) = (15-20)/5 = -1
P(z < -1) = 0.1586..
------------------------------
b. What percent of the invoices are paid in more than 28 days?
z(28) = (28-20)/5 = 1.6
P(z > 1.6) = 0.5480...
------------------------------------
c. What percent of the invoices are paid in more than 15 days
but less than 28 days?
P(-1 < z < 1.6) = 0.7865
-------------------------------------
d. The management of Hub Furniture wants to encourage their customers to pay their monthly invoices as soon as possible. Therefore, it announced that a 2 percent reduction in price would be in effect for customers who pay within 7 days of the receipt of the invoice. What percent of customers will earn this discount?
z(7) = (7-20)/5 = -2.6
P(z < -2.6) = 0.0047
==============================
Cheers,
Stan H.